OpenStudy (anonymous):

how to find the integration of {sin3x/(sin5x*sin2x)}

OpenStudy (experimentx):

try, sin(3x) = sin(5x - 2x) ... then separate fractions!!

OpenStudy (experimentx):

\[ \frac{\sin 5x \; \cos 2x}{ \sin 5x \; \sin 2x} - \frac{\sin 2x \; \cos 5x}{ \sin 5x \; \sin 2x}\] that seems a lot simpler!!